OpenStudy (hy123):

y'' + 4y = 8 y(0)=2, y'(0)=1 Solve the following differential equation using the Laplace transform method.

OpenStudy (hy123):

I know how to take the laplace transform of the left side but im not sure what to do with the 8 on the right side.

OpenStudy (anonymous):

Just do \(8\mathcal L[1]\)

OpenStudy (hy123):

Oh that is right.. So it is basically 8t?

OpenStudy (anonymous):

I guess? Can't remember the transforms.

OpenStudy (hy123):

actually.. 1 = 1/s so therefore it is 8/s

OpenStudy (anonymous):

\(1=t^0\)

OpenStudy (hy123):

I'm looking at m formula sheet and it says 1 = 1/s

OpenStudy (hy123):

Thanks anyways wio!